OpenStudy (anonymous):

Solve the quadratic inequality. Give the answer in interval notation. x^2-7x+16> or equal to 10

OpenStudy (demandandsupply):

rearrange for 0, then factorise

OpenStudy (demandandsupply):

i dont know what interval notation is, but it should give you x is less than or equal to 1, and x is greater than or equal to 6

OpenStudy (cwrw238):

this transposes to x^2 - 7x + 6 > 0 first solve the equation x^2 - 7x + 6 = 0 to get the critical points then you can solve it by considering the values of the function in the intervals of the roots or from the graph of the function

OpenStudy (cwrw238):

you can solve the equation by factoring

OpenStudy (jdoe0001):

\(\bf x^2-7x+16\ge10\implies x^2-7x+16-10\ge0\implies x^2-7x+6\ge 0\\ \quad \\ \textit{quadratic formula}\\ \begin{array}{cccllll} y=&1x^2&-7x&+6\\ &\uparrow&\uparrow&\ \uparrow\\ &a&b&c \end{array}\qquad x= \cfrac{ - b \pm \sqrt { b^2 -4ac}}{2a}\)